OpenStudy (anonymous):

Please help with a few questions? A 175 kg motorcycle has 6.8 x 104J of kinetic energy. How fast is it going?

OpenStudy (anonymous):

@TheSmartOne @Here_to_Help15

OpenStudy (amistre64):

what is the formula that defines kinetic energy?

OpenStudy (anonymous):

Thats all I need the formula

OpenStudy (amistre64):

the question assumes that you have been taught the formula, and is asking you to apply it.

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

what is your best recollection of the formula?

OpenStudy (anonymous):

I can not find it in my formula notes... I have all of them but this one...

OpenStudy (anonymous):

This one and one other

OpenStudy (amistre64):

well, im of the mindset that it would be easier for you to remember/learn if you can find it and post it. try a google search for kinetic energy formula and it should pop up pretty easily. afterwards i can confirm it or correct it if need be.

OpenStudy (amistre64):

even a vague recollection would be useful :)

OpenStudy (anonymous):

k=1/2mv^2

OpenStudy (anonymous):

I have this one in my notes

OpenStudy (amistre64):

very good and all we need to do is solve for v

OpenStudy (anonymous):

175=1/26.8*10^2

OpenStudy (amistre64):

175 kg motorcycle has 6.8 x 104J of kinetic energ \[6.8(10^4)=\frac{175}{2}v^2\] one stray thought, is 175 mass or weight? im assuming its the mass. \[\frac{2}{175}(6.8)(10^4)=v^2\] \[\sqrt{\frac{2}{175}(6.8)(10^4)}=v\]

OpenStudy (anonymous):

so would 10 become 10000?

OpenStudy (amistre64):

sqrt(10^4) = 10^2 = 100

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So it would be 28 m/s

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=sqrt%282%286.8%29%2810%5E4%29%2F175%29 yes, that is what im coming up with as well

OpenStudy (anonymous):

okay I have one more question can you please help me with?

OpenStudy (amistre64):

i can try

OpenStudy (anonymous):

A force of 35 N is used to stretch a spring 15 cm beyond its normal length. What is the increase in the spring’s energy?

OpenStudy (anonymous):

I dont know why there are question marks....

OpenStudy (amistre64):

its an encoding issue on the site. a refresh clears them up

OpenStudy (amistre64):

tends to happen alot for copy/paste stuff

OpenStudy (anonymous):

okay so all my formulas involving this equation need m but I have cm?

OpenStudy (amistre64):

convert if need be 1m = 100cm; therefore 1cm = 1/100 m 15cm = 15/100 m = .15m

OpenStudy (anonymous):

okay thank you for that!

OpenStudy (amistre64):

youre welcome i had to look up an energy equation for this: \[PE=\frac12kx^2\] and using the k from the spring constant:\[F=kx~:~k=F/x\] \[PE=\frac12 Fx\]

OpenStudy (amistre64):

does that seem right to you?

OpenStudy (anonymous):

U=mgh

OpenStudy (anonymous):

I dont know how to fill the numbers in?

OpenStudy (anonymous):

Thats what I found...

OpenStudy (amistre64):

lets use the Fx/2 :) http://www.physicsclassroom.com/Class/energy/u5l1b.cfm

OpenStudy (amistre64):

PE = 35(.15)/2 PE = 35(.075)

OpenStudy (anonymous):

sorry I lost connection @amistre64

OpenStudy (amistre64):

site was on the fritz for me as well.

OpenStudy (anonymous):

I got an extreamly weird answer .07

OpenStudy (amistre64):

i get about 2.6 :)

OpenStudy (amistre64):

from the site i looked at: PE = 1/2 kx^2, given that: F = kx since we have F and x, we can rewrite this as: PE = 1/2 Fx^2/2 or simply: PE = Fx/2

OpenStudy (anonymous):

That makes much more sence!

OpenStudy (amistre64):

good luck